# Exploring Knowledge Boundaries in Large Language Models for Retrieval Judgment

### v1

@@ 4 Can LLM express know/unknown?
In Section §3.2, we noted that using confidence and certainty as screening indicators helps identify questions with high mastery and low confidence for retrieval enhancement. This approach can reduce the proportion of retrieval and provide performance benefits. So, how can we allow LLMs to express their mastery and certainty? We explore two simple approaches to examine two key questions: Can the model effectively express mastery and certainty? Are there internal model features or data distribution characteristics related to knowledge that the model does not master or finds uncertain?
@@ 4.1 Methods
To construct training data for the KBM, we generate soft labels using the sampling method detailed in Section §3.2. As shown in Figure 4, each query is assessed under two configurations: LLM Only and RAG. Based on accuracy and certainty, we derive two data fields: mastery and certainty. We establish a threshold, $\tau$, to differentiate between known and unknown data, with values below $\tau$ labeled as False and those above $\tau$ labeled as True. In the training phase, the generated data is formatted into question-answer pairs and utilized to fine-tune the sampled ontology model (Qwen2-7B instruct). Further details regarding the settings can be found in the Appendix §A.1.
@@
During the inference phase of the KBM, the process unfolds in three steps: At step 1, after the user enters a query, the system packages this query through a prompt and sends it to the KBM to assess mastery and certainty. At step 2, if the KBM judges the query as True, it forwards the original query directly to the answer generation model (Qwen2-7B instruct) to generate a response. If the query is judged as False, the process proceeds to step 3. At step 3, the system performs an open-domain retrieval using Google, based on the user’s original query. The retrieved chunks of information are then spliced together. The original query and the connected chunks are combined using RAG Prompt and subsequently fed into the answer generation model to produce the final response.
4.2 BASELINES AND DATASET
Baselines and Metrics. We establish a baseline using the following methods: (1) Prompting Methodology: We utilize state-of-the-art models, including LLAMA3-70B, Qwen2-72B, and GPT-4. Each model is presented with the same retrieval query. If a model indicates a need for additional information, we perform a Google search and enhance the response using the top 10 results through RAG. If no retrieval is necessary, the model generates the answer autonomously. (2) Random: This method serves as a dynamic benchmark, involving random extractions of queries for RAG. The search proportions align with the retrieval ratios established by each baseline method. Each methodology offers a unique approach to managing knowledge boundaries, contributing to a comprehensive evaluation of our proposed frameworks. In terms of performance Metrics, we use Exact Match (EM) for NaturalQA and TrivialQA, and Accuracy for the other datasets, while also considering the retrieval ratio (Rat.) as an important metric. Further details can be found in the Appendix §A.2.
Dataset. NaturalQA (Kwiatkowski et al., 2019): This QA dataset, curated by Google, consists of real-world questions derived from natural retrieval queries. TriviaQA (Joshi et al., 2017): This dataset is based on encyclopedic content and features complex questions and answers, primarily sourced from competitions and quizzes. MMLU (Hendrycks et al., 2021): This dataset comprises multiple-choice questions across various fields, assessing the model’s knowledge mastery in academic and professional domains. OpenBookQA (Mihaylov et al., 2018): Focusing on scientific inquiries, this dataset requires reasoning rooted in principles and common sense. en-SQuAD-en2.0 (Rajpurkar et al., 2018): This dataset features question-answer pairs and evaluates reading comprehension skills. FreshQA-en (Vu et al., 2023): This dataset presents various question and answer types, offering a comprehensive assessment of QA capabilities. HotpotQA (Yang et al., 2018): This dataset consists of 113,000 Wikipedia based QA pairs that necessitate complex reasoning across multiple supporting documents and include sentence-level supporting facts. For the Chinese dataset, we utilize the following resources: WebQA (Chang et al., 2022): This open-domain QA dataset is collected via web crawlers, covering a wide array of topics and evaluating the performance of QA systems. SogouQA  [^3] https://github.com/sherlcok314159/ChineseMRC-Data: Provided by Sogou, this dataset features user-generated questions and system-generated answers, assessing accuracy and robustness. MLEC (Li et al., 2021): This dataset is designed to test the comprehension capabilities of models in various contexts. Xiezhi (Gu et al., 2024): A set of 249,587 Chinese/English questions covering 516 subjects for evaluating LLMs. SQuAD-zh  [^4] https://github.com/pluto-junzeng/ChineseSquad: The Chinese version of the English SQuAD (Rajpurkar et al., 2018) dataset serves to train and evaluate machine reading comprehension and QA systems. C-EVAL (Huang et al., 2023): This comprehensive Chinese evaluation suite assesses large language models’ knowledge and reasoning across 52 disciplines and four difficulty levels. CMMLU (Li et al., 2024a): This comprehensive benchmark assesses the knowledge and reasoning capabilities of language models across 67 topics, from basic to advanced. We used WebQA, TriviaQA, and a combination of MMLU, MLEC, and XieZhi training sets to train the KBM model. The test set and specific data information are shown in the Appendix.
5 Simulate knowledge boundary
Mastery Certainty Prompt LLM All
A/E Rat(%) Rnd A/E Rat(%) Rnd A/E Rat(%) Rnd Only RAG
KBM (Qwen2-7B-Instruct based)
en-NaturalQA 60.09 86.15% 51.20 61.27 97.53% 55.16 59.92 85.98% 57.91 35.58 61.63
en-TirviaQA 81.39 70.25% 74.42 83.32 87.70% 80.66 79.08 73.55% 75.62 50.79 84.60
en-OpenBoook 82.64 26.20% 81.92 82.32 57.80% 82.88 81.96 27.80% 82.12 82.72 81.48
en-MMLU 70.00 70.01% 68.59 70.45 65.31% 69.14 67.30 27.24% 67.70 66.45 70.80
en-SQuAD 89.13 87.18% 83.14 89.38 97.27% 88.61 78.55 75.55% 75.30 34.02 89.83
en-FreshQA 58.40 86.80% 57.80 59.72 91.80% 58.56 58.52 86.80% 57.82 33.68 60.84
[…]
zh-SogouQA 81.59 75.23% 74.17 82.89 94.56% 81.06 78.13 78.90% 76.75 52.53 83.01
zh-SQuAD 82.66 96.00% 81.29 83.45 99.33% 83.31 78.74 90.33% 77.82 22.72 83.66
zh-CMMLU 78.32 49.98% 77.73 77.12 95.29% 77.06 77.97 49.71% 77.87 78.35 77.05
zh-Ceval 73.56 52.11% 71.99 71.71 94.03% 71.69 72.89 51.28% 72.39 72.93 71.62
Table 1: Evaluation results of KBM on 11 test sets. LLM Only refers to baselines operating without retrieval, while ALL RAG represents baselines utilizing RAG.
@@ 5 Simulate knowledge boundary
Figure 5: The performance of LLM in the case of KBM model being judged as Known/Unknown.
@@ 5.1 End-to-end evaluation
We evaluate the end-to-end performance of KBM across 11 test sets, demonstrating that both Mastery and Certainty reduce the search ratio while enhancing the LLM’s ability to answer questions. Mastery adopts a more conservative approach to search ratio management, while Certainty emphasizes performance improvement. The results are presented in Table 1.
Our analysis reveals that for short answer and QA test sets, such as en-NaturalQA, en-TriviaQA, zh-WebQA, and zh-SogouQA, both methods outperform the random baseline, with Mastery achieving particularly notable results. In reading comprehension tasks like en-SQuAD and zh-SQuAD, where background information is crucial for accurate answers, KBM exhibits a higher search ratio. All methods demonstrate strong performance and a high search ratio on en-FreshQA, indicating that KBM and LLM effectively capture queries with temporal features. However, in the context of multiple-choice questions found in CMMLU, MMLU, and OpenBookQA, our method shows only marginal improvements over the random baseline. We suspect this limited enhancement arises from the challenges in sourcing relevant information for multiple-choice formats.
KBM enables the model to effectively distinguish between known and unknown information, thereby optimizing performance and reducing costs. We analyze the mean scores across all test sets when categorized as known or unknown, as illustrated in Figure 6. The results indicate that when classified as known, the mean answer score of the LLM is higher, while the score decreases when information is deemed unknown. This pattern aligns with our expectations regarding the differentiation process. In contrast, although the prompt-based method can distinguish between known and unknown information, its effectiveness is comparatively lower.
@@
KBM
-Mastery 94.6% 51.6 92.3% 59.0 75.9% 62.7
-Certainty 98.5% 52.0 95.6% 60.2 89.4% 64.6
@@
Table 2: Retrieval judgment results of dynamic changing class tests.
@@
Table 3: Retrieval judgment results of multi-hop class tests.
5.2 Analysis
Dynamic Knowledge: Utilizing KBM for Identifying Questions with Temporal Variability. We demonstrate that KBM effectively identifies questions with answers that change over time. Specifically, we classify the temporal changes in answers found in the FreshQA dataset into three distinct categories: fast-changing, slow-changing, and never-changing, based on the frequency of these changes. In open domains, variations in answers often signal the need for knowledge updates, necessitating the integration of external information into the LLM. As illustrated in Table 3, the end-to-end performance without RAG for both fast-changing and slow-changing categories is suboptimal, indicating a reliance on external knowledge. The high retrieval rates of KBM for both fast-changing and slow-changing categories suggest that it adeptly captures these evolving answers. Conversely, the lower retrieval rate for the never-changing category implies that some knowledge is effectively embedded within the LLM. This observation underscores KBM’s sensitivity in identifying questions with temporally correlated answers, highlighting its role in enhancing dynamic knowledge adaptation.
Long-tail Static Knowledge: Evaluating KBM’s Performance Across Knowledge Frequencies. We investigate the capacity of KBM to capture low-frequency long-tail knowledge across various question sets. Specifically, we combine test data from WebQA, SogouQA, and SQuAD-zh. Utilizing the gold answers from these datasets, we conduct vector retrieval within our Chinese database to differentiate knowledge based on its frequency. As illustrated in Figure 6, the LLM Only approach demonstrates reduced accuracy for low-frequency knowledge answers while performing better for high-frequency knowledge. However, integrating KBM with a prompt-based retrieval mechanism significantly enhances the model’s performance for long-tail low-frequency knowledge. Notably, the Certainty-based method yields the most substantial improvement, followed by the Mastery-based approach. These findings indicate that KBM effectively detects low-frequency long-tail knowledge and boosts overall performance through strategic retrieval.
@@ 5.2 Analysis
Multi-Hop: Multi-hop Knowledge Detection with KBM. A crucial aspect of our analysis is the ability of KBM to detect complex queries that necessitate multi-hop knowledge. Multi-hop questions comprise intricate knowledge components, requiring adjustments to the LLM. In these scenarios, KBM identifies the complexity of the queries and effectively employs RAG. We tested queries involving two or more hops from the FreshQA and Hotpot test sets, with results presented in Table 6. KBM demonstrates higher retrieval rates for multi-hop questions, indicating its effectiveness in navigating these complexities. While the overall end-to-end improvement is modest, this limitation stems from the need for further optimization of the RAG pipeline.
KBM-Mastery KBM-Certainty Prompt LLM Only ALL RAG
A/E Rat(%) Rnd A/E Rat(%) Rnd A/E Rat(%) Rnd A/E A/E
KBM For zh-WebQA
-GPT-4 85.7 57% 78.5 89.5 90% 87.8 69.5 14% 68.0 64.3 90.0
-Qwen2-72B 91.4 57% 88.2 91.8 90% 91.3 87.9 22% 85.6 83.7 91.9
-Llama3-70B 77.9 57% 72.6 88.2 90% 86.7 59.7 16% 55.3 48.7 90.9
[…]
-GPT-4o 60.6 86% 60.4 60.7 98% 60.5 55.7 18% 56.6 55.6 60.1
-Qwen2-72B 58.5 86% 56.5 59.3 98% 58.9 54.7 51% 52.6 42.7 59.4
-Llama3-70B 59.6 86% 58.6 60.2 98% 59.7 59.2 47% 58.6 47.2 60.2
@@
Figure 7: Illustration of KBM exploring the boundaries of RAG.
Figure 8: Comparison of KBM double judgment results under RAG.

### v2

@@ 3 Can LLM Express Know/Unknown?
In this section, we generate known and unknown soft labels based on accuracy and certainty for training the KBM, and we introduce the baselines and datasets used for evaluation.
NQ TriQA OBQA MMLU SQD MLEC WebQA SogQA zh_SQD CMMLU CSQA Avg.
Naive (Qwen2-7B-Instruct based + Google)
LLM A/E 35.58 50.79 82.72 66.45 34.02 74.25 70.06 52.53 22.72 74.23 24.13 53.41
RAG A/E 61.63 84.60 81.48 70.80 89.83 68.40 92.54 83.01 83.66 77.05 77.80 79.16
Qwen2-7B-Instruct based + Google
Prompt A/E 59.92 79.08 81.96 67.30 78.55 72.80 89.75 78.13 78.74 77.97 75.41 76.33
Rnd. 57.91 75.62 82.12 67.70 75.30 72.01 87.01 76.75 77.82 72.15 75.14 74.50
Rat. 86.0% 73.6% 27.8% 27.2% 75.6% 44.7% 77.5% 78.9% 90.3% 49.7% 94.1% 65.9%
Self-RAG (Llama-2 based + Google)
 - 7B A/E 31.07 69.82 74.80 47.93 62.92 29.48 52.96 39.83 29.47 33.94 18.03 44.57
Rnd. 35.79 69.49 76.40 49.38 62.27 30.72 54.02 39.83 30.33 34.63 19.17 45.64
Rat. 22.7% 55.9% 10.8% 29.9% 77.0% 10.4% 28.3% 28.5% 37.0% 3.4% 23.2% 29.7%
KBM (Qwen2-7B-Instruct based + Google)
 - Acc. A/E 60.09 81.39 82.64 70.00 89.13 73.28 90.84 81.59 82.66 78.32 76.57 78.77
Rnd. 51.20 74.42 81.92 68.59 83.14 72.66 81.19 74.17 81.29 77.73 74.87 74.65
Rat. 86.2% 70.3% 26.2% 70.0% 87.2% 30.2% 56.8% 75.2% 96.0% 50.0% 94.3% 67.5%
 - Cer. A/E 61.27 83.32 82.32 70.45 89.38 71.69 92.12 82.89 83.45 77.12 77.38 79.22
Rnd. 55.16 80.66 82.88 69.14 88.61 69.46 90.71 81.06 83.31 77.06 76.85 77.72
Rat. 97.5% 87.7% 57.8% 65.3% 97.3% 69.0% 89.7% 94.6% 99.3% 95.3% 97.8% 86.5%
Table 1: Comparison of performance and trigger ratio (Rat. %) metrics between KBM and baseline models. The English dataset utilizes EM as the performance metric, while the Chinese dataset employs accuracy as the evaluation metric. Each method is compared with the random RAG trigger score (Rnd.) at the same trigger ratio.
@@ 3.1 Methods
To construct training data for the KBM, we generate soft labels using the sampling method detailed in Section §2.2. As shown in Figure 3, each query is assessed under two configurations: LLM Only and LLM + RAG, providing scores for accuracy and certainty. A threshold of $\tau=0.9$ is set to classify data as known or unknown. The data is then formatted into QA pairs for fine-tuning the Qwen2-7B model. For additional setup details, refer to Appendix §B.
@@
During the inference phase of the KBM, the process unfolds in three steps: At step 1, after the user enters a query, the system packages this query through a prompt and sends it to the KBM to assess accuracy and certainty. At step 2, if the KBM judges the query as known, it forwards the original query directly to the answer generation model (Qwen2-7B instruct) to generate a response. If the query is judged as unknown, the process proceeds to step 3. At step 3, the system performs an open-domain retrieval using Google, based on the user’s original query. The retrieved chunks of information are then spliced together. The original query and the connected chunks are combined using RAG prompt and subsequently fed into the answer generation model to produce the final response.
3.2 Experiment Setting
Metrics: For the English test data, we use Exact Match (EM) as the metric, while Accuracy is applied to the other datasets, represented as A/E. Additionally, we consider the retrieval ratio (Rat.) as a crucial metric.
Baselines: We establish a baseline using the following methods: (1) Prompt: we use prompts from Qwen2-7B Yang et al. (2024), Llama3-70B AI@Meta (2024), Qwen2-72B Yang et al. (2024), and GPT-4o OpenAI et al. (2024) to determine whether the current query requires retrieval. Each model is presented with the same retrieval query. (2) Random: This method serves as a dynamic benchmark. By analyzing the ratio of each baseline method that triggers RAG, a query from the test set is randomly selected to initiate the RAG evaluation. This approach effectively assesses whether the model has genuinely identified the unknown query or merely "guessed" it. Among related approaches, we select Self-RAG (Llama-2 7B based) (Asai et al., 2023) as baselines for our analysis. We use Google as the retriever for all methods, selecting the first 10 snippets as contextual information. The experiment focuses on end-to-end effectiveness, the reduction in retrieval ratio, and the differences between the baseline methods and the random approach.
Datasets: KBM’s training and test data encompass a variety of task types, including short answer questions, multiple choice questions, reading comprehension, and multi-hop questions. We utilize TriviaQA (TriQA) Joshi et al. (2017), WebQA Chang et al. (2022), and a combination of MMLU Hendrycks et al. (2021), MLEC Li et al. (2021), and XieZhi Gu et al. (2024) training sets to train the KBM model. We categorize the test sets into two groups: In-Domain, which includes TriviaQA Joshi et al. (2017), WebQA Chang et al. (2022), MLEC (Li et al., 2021), and MMLU Hendrycks et al. (2021), and Out-Of-Domain, which comprises NaturalQA(NQ) Kwiatkowski et al. (2019), Open BookQA (OBQA) Mihaylov et al. (2018), SQuAD (SQD) Rajpurkar et al. (2018), FreshQA Vu et al. (2023), SogouQA (SogQA)  [^3] https://github.com/sherlcok314159/ChineseMRC-Data, the Chinese SQuAD (zh_SQD)  [^4] https://github.com/pluto-junzeng/ChineseSquad, CMMLU Li et al. (2024a), and Chinese SimpleQA (CSQA) He et al. (2024). For details on the datasets, see the Appendix §B.2.
@@ 3.2 Experiment Setting
Figure 4: The performance of LLM in the case of KBM model being judged as known/unknown (0-100). The higher the average score of known, the better.
Figure 5: Illustration of the KBM method’s effects: the top shows a decrease in harmful counts, while the bottom indicates an increase in beneficial counts.
4 Experiments
In this section, we design extensive experiments to verify the effectiveness of KBM.
@@ 4.1 End-to-End Evaluation
We evaluate the end-to-end performance of KBM across 11 test sets, showing that both accuracy and certainty reduce the RAG trigger ratio while enhancing the LLM’s ability to answer questions. The results are presented in Table 1. On average, the KBM accuracy-based method outperformed the LLM Only method by 25.4%, narrowing the gap with All RAG to 0.4%, while decreasing the RAG triggering ratio by 32.5%, which is 4.1% higher than the Random method, indicating a greater beneficial density. Similarly, the KBM certainty-based method improved performance by 25.8% compared to the LLM Only method, achieving a performance that is 0.1% higher than All RAG, and reduced the retrieval ratio by 13.5%. Compared with the Prompt and Self-RAG methods, KBM effectively reduces the RAG triggering proportion while maintaining high performance.
Specifically, our analysis reveals that for short answer and QA test sets (e.g., NQ, TriviaQA, WebQA, SogouQA, and CSQA), both methods outperform the random baseline, with accuracy achieving particularly notable results. In reading comprehension tasks like SQA and zh_SQD, where background information is essential for accurate answers, KBM exhibits a higher trigger ratio. All methods demonstrate strong performance and a high trigger ratio on FreshQA, indicating that KBM and LLM effectively capture queries with temporal features. However, in the context of multiple-choice questions found in CMMLU, MMLU, and OBQA, our method shows only marginal improvements over the random baseline. We guess that this limited enhancement arises from the challenges of sourcing relevant information for multiple-choice formats. Moreover, the diversity of training data formats, in relation to data type and task domain, is crucial for the effectiveness of KBM.
Known/Unknown Expression Analysis. KBM enables the model to effectively distinguish between known and unknown information, thereby optimizing performance and reducing costs. We analyze the mean scores across all test sets categorized as known or unknown by KBM, as illustrated in Figure 4. We can see that without the assistance of RAG (represented by the orange bar), the LLMs struggle to handle unknown questions. In contrast, the two KBM methods achieve the most significant improvements compared to other approaches. Specifically, the prompt-based method shows ‘confusion’ about the knowledge boundary because there is a notable gap between Known Aeverage Scores and Unknown Average Scores with RAG. For improvements in accuracy and certainty at finer intervals, see the Appendix §A.1.
@@
Naive (Qwen2-7B-Instruct based + Google)
LLM 0% 29.7 0% 29.8 0% 40.2
RAG 100% 50.4 100% 60.2 100% 66.6
@@
Self-RAG(7B) 56.2% 37.4 56.3% 49.4 55.1% 56.0
KBM-Acc. 94.6% 51.6 92.3% 59.0 75.9% 62.7
KBM-Cer. 98.5% 52.0 95.6% 60.2 89.4% 64.6
Table 2: Results of triggering RAG with various levels of dynamic knowledge.
Impact Analysis of KBM on RAG. In Section §2 Figure 2, we discuss the three impacts of RAG on LLM. To investigate the changes in the number of harmful and beneficial questions, we employ KBM instead of RAG and find that KBM reduces the number of harmful questions while increasing the number of beneficial ones. Specifically, as shown in Figures 5(a) and 5(b), using WebQA as an example, we find that compared with RAG, the KBM-based approach decreases the number of harmful questions across all five intervals. This reduction is particularly evident in the high accuracy interval, indicating that we mitigate the harmful impact associated with RAG in areas close to what is known. Similarly, as illustrated in Figures 5(c) and 5(d), KBM increases the number of beneficial questions in each interval, showing that it enhances system performance by effectively triggering RAG. Moreover, we conduct a comprehensive efficiency analysis, the results of which are presented in §A.3.
4.2 Analysis of Complex Scenarios
A multitude of complex scenarios necessitate the frequent invocation of RAG. In light of this, we conducted tests and focused on the ratio of KBM in the following scenarios.
Dynamic Knowledge. We demonstrate that KBM effectively identifies questions with answers that change over time. Specifically, we classify the temporal changes in answers found in the FreshQA dataset into three distinct categories: fast-changing, slow-changing, and never-changing, based on the frequency of these changes. In open domains, variations in answers often indicate the need for knowledge updates, necessitating the integration of external information into the LLM. As shown in Table 2, the end-to-end performance without RAG for both fast-changing and slow-changing categories is suboptimal, highlighting a reliance on external knowledge. The high RAG trigger rates of KBM for these categories suggest that it effectively captures evolving answers. Conversely, the lower retrieval rate for the never-changing category implies that some knowledge is effectively embedded within the LLM. This finding underscores KBM’s sensitivity in identifying questions with temporally correlated answers and highlights its role in enhancing dynamic knowledge adaptation.
@@ 4.2 Analysis of Complex Scenarios
Self-RAG(7B) 64.3% 47.5 65.0% 38.6
KBM-Acc. 96.5% 48.9 91.2% 50.9
KBM-Cer. 99.1% 48.7 94.3% 51.4
Table 3: Results of multi-hop triggering RAG.
@@
Multi-Hop. A crucial aspect of our analysis is the ability of KBM to detect complex queries that necessitate multi-hop knowledge. Multi-hop questions comprise intricate knowledge components, requiring adjustments to the LLM. In these scenarios, KBM identifies the complexity of the queries and effectively employs RAG. We tested queries involving two or more hops from the FreshQA($\geq$2-hop) and HotpotQA test sets, with results presented in Table 3. KBM demonstrates higher retrieval rates for multi-hop questions, indicating that it reveals more unknowns when addressing these complexities. As a result, the trigger ratio is higher than the average ratio for the task. While the overall end-to-end improvement is modest, this limitation stems from the need for further optimization of the RAG pipeline. We employ KBM to assess its ability to detect complex problems and trigger RAG, rather than to break them down and resolve them.
Long-tail Static Knowledge. Long-tail knowledge has consistently posed challenges for the learning process of LLMs Kandpal et al. (2023). We investigate the capacity of KBM to capture low-frequency long-tail knowledge across various question sets, confirming its effectiveness across all frequency ranges. Specifically, we combine test data from WebQA, SogQA, and zh_SQD. Utilizing the gold answers from these datasets, we conduct vector retrieval within our Chinese database to differentiate knowledge based on its frequency. As illustrated in Figure 6, the LLM Only approach demonstrates reduced accuracy for low-frequency knowledge answers while performing better for high-frequency knowledge. However, integrating KBM with a prompt-based retrieval mechanism significantly enhances the model’s performance for long-tail low-frequency knowledge. Notably, the certainty-based method yields the most substantial improvement, followed by the accuracy-based approach. These findings indicate that KBM effectively detects low-frequency long-tail knowledge and boosts overall performance by triggering RAG.
  Gpt-4o   Qwen2-72B Llama3-70B
WebQA NQ WebQA NQ WebQA NQ
LLM A/E 64.3 55.6 83.7 42.7 48.7 47.2
RAG A/E 90.0 60.1 91.9 59.4 90.9 60.2
Prompt A/E 69.5 55.7 87.9 54.7 59.7 59.2
Rnd. 60.8 56.6 85.6 52.6 55.3 58.6
Rat. 14% 18% 22% 51% 16% 47%
Self-RAG
A/E 72.6 56.7 86.5 45.7 60.9 49.6
Rnd. 71.6 57.1 86.0 42.7 48.7 47.2
Rat. 28% 23% 28% 23% 28% 23%
KBM-Acc. A/E 85.7 60.6 91.4 58.5 77.9 59.6
Rnd. 78.5 60.4 88.2 56.5 72.6 58.6
Rat. 57% 86% 57% 86% 57% 86%
KBM-Cer. A/E 89.5 60.7 91.8 59.3 88.2 60.2
Rnd. 87.8 60.5 91.3 58.9 86.7 59.7
Rat. 90% 98% 90% 98% 90% 98%